OpenStudy (anonymous):

A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound? A. 15 lbs oatmeal, 15 lbs bran flakes B. 20 lbs oatmeal, 10 lbs bran flakes C. 15 lbs bran flakes, 15 lbs oatmeal D. 20 lbs bran flakes, 10 lbs oatmeal

OpenStudy (anonymous):

3.60x+4.80y=4(30)

OpenStudy (anonymous):

? i dont understant explain please

OpenStudy (anonymous):

im sorry i meant to include an explaination. so you know that you have to sell a certain number of pounds of bran, and a certain number of pounds of oatmeal. but you arent sure what those pounds are, but you know that they have to equal $4.00 a pound, and you want to sell 30 pounds of it. so you are going to multiply 30lbs by $4.00 to get the total money earned from it, and set it equal to the lbs of bran plus lbs of oatmeal, and what they sell for. do i need to clarify further?

OpenStudy (anonymous):

okay so after i multiply then what yes clarify

jimthompson5910 (jim_thompson5910):

let x = # of pounds of oatmeal y = # of pounds of bran flakes you want to make a 30 lb mix, so x+y = 30 ---> y = 30 - x ------------------------------------ We know that "A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound" and we also know that want the 30 lb mix to be sold at $4 a pound, so 3.60x + 4.80y = 30*4 3.60x + 4.80y = 120 3.60x + 4.80(30 - x) = 120 ... plug in y = 30 - x 3.60x + 144 - 4.80x = 120 I'll let you finish up

OpenStudy (anonymous):

so its B!

jimthompson5910 (jim_thompson5910):

yes, nice work